# Q-Probe: A Lightweight Approach to Reward Maximization for Language Models

### v1

@@ 6.3 OpenAI API results
Finally, we conduct a similar experiment on top of generations from the OpenAI API. Results are reported in Table 2. We use embeddings from CodeLlama-70b-Python, since embeddings are not available from the API generative model. We find gains over the baselines on both datasets.
@@ 7 Preference feedback experiments
Table 3: Comparison of different preference learning methods on the combination of three datasets. [Ethayarajh et al., 2023]’s setting is exactly followed; numbers for base models are taken from their paper. Base model is LLaMA-7B after SFT training.
@@
7.1 Setup
We first extract features for probe training. Combining the training sets of three datasets together, we obtain a dataset with $200,336$ training pairs, each containing a winning completion and a losing completion. We concatenate the prompt with both completions and run a forward pass of the model to extract embeddings. Note that our Q-probing is applied on the supervised finetuned model, which is also the starting point for the compared methods [Ouyang et al., 2022, Rafailov et al., 2023, Ethayarajh et al., 2023]. Offline PPO, DPO and KTO use different loss functions to finetune the model weights from this supervised finetuned model.
Experiment Details
7.2 Experimental Results
Figure 5: How the win rate of Q-probe scales with inference-time compute on preference learning benchmarks. The skyline shows the performance of a perfect oracle selector. Shade area represents $95\%$ confidence interval for $10$ runs.
@@ 8 Discussion
We have proposed Q-probe, a lightweight replacement for or complement to finetuning when we want to maximize reward on downstream tasks given a pre-trained language model. On two settings with access to oracle rewards and human preference pairs respectively, Q-probe outperforms strong finetuning baselines. For anyone who does not have the resource or access to finetune large language models but wish to adapt them for their own downstream tasks, Q-probe can serve as a solid replacement, and even given a finetuned model, Q-probe can be added on top to leverage more inference-time compute to squeeze out better performance.

### v2

@@ 6.3 OpenAI API results
Finally, we conduct a similar experiment on top of generations of the OpenAI API. Results are reported in Table 2. We use embeddings from CodeLlama-70b-Python since embeddings are not available from the API generative model. We find gains over the baselines on both datasets.
@@
6.4 Additional Experiments on GSM-8K
We also conduct experiment on GSM-8K with Code-Llama-7B, $k=48$ and $\beta=0.1$, following the implementation of [Gao et al., 2022, Cobbe et al., 2021], using 8-shot evaluation with code adopted from the Code Generation LM Evaluation Harness project [Ben Allal et al., 2022]. Results in Table 3 show a similar trend as the experiments on coding.
Table 3: Expected return on GSM-8K, trained on 7473 problems from the training set of GSM-8K. Hyperparameters kept the same as Table 1. Evaluation protocols follow [Gao et al., 2022].
Method GSM-8k
Baseline (Pass@1) 0.25
Baseline (Greedy) 0.29
Q-probe $L_{Q}$ 0.36
Q-probe $L_{CE}$ 0.43
Q-probe $L_{PG}$ 0.45
(Skyline) Pass@48 0.80
@@ 7 Preference feedback experiments
7.1 Setup
We first extract features for probe training. Combining the training sets of three datasets together, we obtain a dataset with $200,336$ training pairs, each containing a winning completion and a losing completion. We concatenate the prompt with both completions and run a forward pass of the model to extract embeddings. Note that our Q-probing is applied on the supervised finetuned model, which is also the starting point for the compared methods [Ouyang et al., 2022, Rafailov et al., 2023, Ethayarajh et al., 2023]. Offline PPO, DPO, and KTO use different loss functions to finetune the model weights from this supervised finetuned model.
Experiment Details
7.2 Experimental Results
Figure 5: How the win rate of Q-probe scales with inference-time compute on preference learning benchmarks. The skyline shows the performance of a perfect oracle selector. The shaded area represents $95\%$ confidence interval for $10$ runs.
Table 4: Comparison of different preference learning methods on the combination of three datasets. [Ethayarajh et al., 2023]’s setting is exactly followed; numbers for base models are taken from their paper. The base model is LLaMA-7B after SFT training.
@@ 8 Discussion
We have proposed Q-probe, a lightweight approach to maximize reward on downstream tasks given a pre-trained language model. Q-probes can be used effectively as a complement to or replacement for other techniques like finetuning or prompting. On two settings with access to oracle rewards and human preference pairs respectively, Q-probe outperforms strong baselines. For anyone who does not have the resource or access to finetune large language models but wishes to adapt them for their own downstream tasks, Q-probe can serve as a solid replacement, and even given a finetuned model, Q-probe can be added on top to leverage more inference-time compute to squeeze out better performance.